What is -x over 4 =18 2. What is -2 over 3x = -1 over six
3. What is 14y =23y-17-10 4. What is 0.22y-0.6=0.12y+3-0.8y Please help!

Answer by checkley71(8403)

-X/4=18/1 CROSS MULTIPLYING WE GET -X=18*4 OR -X=72 OR X=-72
-2/3X=-1/6 CROSSMULTIPLYING WE GET -12=-3X OR X=-12/-3 OR X=4
14Y=23Y-17-10 OR 14Y-23Y=-27 OR -9Y=-27 OR Y=-27/-9 OR Y=3
.22Y-0.6=.12Y+3-0.8Y OR .22Y-.12Y+.8Y=3+.6 OR .18Y=3.6 OR Y=3.6/.18 OR Y=20